OpenStudy (anonymous):

The sum of a student's three exam scores is 230. If the first is 10 points more than the second, and the sum of the first two is 14 more than twice the third, what was the first score.

OpenStudy (anonymous):

Let a, b and c be the first, second and the third numbers respectively. From the problem text: a=b+10, a+b=2 c+14, a+b+c=230 Solve the above for a, b and c. {a = 84, b = 74, c = 72}